# Nested subspace learning with flags

### v1

@@ Abstract
Many machine learning methods look for low-dimensional representations of the data. The underlying subspace can be estimated by first choosing a dimension $q$ and then optimizing a certain objective function over the space of $q$-dimensional subspaces—the Grassmannian. Trying different $q$ yields in general non-nested subspaces, which raises an important issue of consistency between the data representations. In this paper, we propose a simple trick to enforce nestedness in subspace learning methods. It consists in lifting Grassmannian optimization problems to flag manifolds—the space of nested subspaces of increasing dimension—via nested projectors. We apply the flag trick to several classical machine learning methods and show that it successfully addresses the nestedness issue.
Keywords: subspace learning, Grassmann manifolds, flag manifolds, nested subspaces
@@ 1 Introduction
Finding low-dimensional representations of datasets is a quite common objective in machine learning, notably in virtue of the curse of dimensionality (Bellman, 1984). A classical way of building such a representation is by searching for a low-dimensional subspace that well represents the data, where “well” is defined by an application-dependent criterion (Cunningham and Ghahramani, 2015). Here is a non-exhaustive list of examples of subspace learning problems. Principal component analysis (PCA) (Jolliffe, 2002) searches for a low-dimensional subspace that minimizes the average squared Euclidean distance to the data. Robust subspace recovery (RSR) (Lerman and Maunu, 2018b) minimizes the average (absolute) Euclidean distance to the data, which is less sensitive to outliers. Many matrix decomposition methods, like robust PCA (Candès et al., 2011) and matrix completion (Keshavan et al., 2010; Candès and Recht, 2012) look for low-rank approximations of the data matrix, which can be decomposed into a product of subspace and coordinate matrices. Trace ratio (TR) (Ngo et al., 2012) refers to a wide class of problems that look for subspaces making a tradeoff between desired yet antagonist properties of low-dimensional embeddings—for instance Fisher’s linear discriminant analysis (LDA) (Fisher, 1936) which seeks to maximize the between-class variance while minimizing the within-class variance. Domain adaptation methods learn some domain-invariant subspaces (Baktashmotlagh et al., 2013) by minimizing the projected maximum mean discrepancy (MMD) (Gretton et al., 2012) between the source and target distributions. Subspace tracking methods incrementally minimize a distance between the current subspace and the available data (Balzano et al., 2010). Subspaces can also be estimated not from the data but from their adjacency matrix or graph Laplacian, as done in the celebrated Laplacian eigenmaps (Belkin and Niyogi, 2003) and spectral clustering (Ng et al., 2001). Subspaces can also be estimated beyond Euclidean data, for instance on symmetric positive definite (SPD) matrices datasets (Harandi et al., 2018). In all the previously cited examples, the search space—the space of all linear subspaces of dimension $q$ embedded in an ambient space of larger dimension $p$—is called the Grassmannian, or the Grassmann manifold (Bendokat et al., 2024) and is denoted $\operatorname{Gr}(p,q)$. Consequently, many machine learning methods can be recast as an optimization problem on Grassmann manifolds.111 Some alternative methods to subspace learning via Grassmannian optimization exist. Notably, a large family of methods—like partial least squares (PLS) (Geladi and Kowalski, 1986), independent component analysis (ICA) (Hyvärinen and Oja, 2000), canonical correlation analysis (CCA) (Hardoon et al., 2004) or projection pursuit (Huber, 1985)—builds a low-dimensional subspace out of the data sequentially. They first look for the best 1D approximation and then recursively add dimensions one-by-one via deflation schemes or orthogonality constraints (Lerman and Maunu, 2018b, Section II.C). Those sequential methods however suffer from many issues—from a subspace estimation point of view—due to their greedy nature. Those limitations are well described in Lerman and Maunu (2018b, Section III.B) and can be extended beyond robust subspace recovery. In light of the fundamental differences between those methods and the previously described subspace methods, we decide not to address sequential methods in this paper.
@@
Learning a low-dimensional subspace from the data through Grassmannian optimization requires to choose the dimensionality $q$ a priori. This prerequisite has many important limitations, not only since a wrong choice of dimension might remove all the theoretical guarantees of the method (Lerman and Maunu, 2018b), but also because the observed dataset might simply not have a favored dimension, above which the added information is invaluable. [^2] Many generative models assume that the data lies on a lower dimensional subspace, up to isotropic Gaussian noise (Tipping and Bishop, 1999; Lerman et al., 2015). However, anisotropy is much more realistic (Maunu et al., 2019), which nuances the theoretical guarantees and removes hopes for a clear frontier between signal and noise at the intrinsic dimension. For these reasons, one might be tempted to “try” several dimensions—i.e. run the subspace optimization algorithm with different dimensions and choose the best one a posteriori, with cross-validation or statistical model selection techniques for instance. In addition to being potentially costly, such a heuristic raises an important issue that we later refer to as the nestedness issue: the optimal subspaces at different dimensions are not nested in general. This notably means that the data embeddings at two different dimensions might drastically differ, which is a pitfall for data analysis. For more details about the importance of nestedness in statistics, one can refer to Huckemann et al. (2010); Jung et al. (2012); Damon and Marron (2014); Huckemann and Eltzner (2018); Pennec (2018); Lerman and Maunu (2018b); Dryden et al. (2019); Yang and Vemuri (2021); Fan et al. (2022). We illustrate in Figure 1 the nestedness issue on toy datasets related to three important machine learning problems: robust subspace recovery, linear discriminant analysis and sparse spectral clustering (Lu et al., 2016; Wang et al., 2017). [^3] The methodology will be provided in due time.
@@
In this paper, we propose a natural solution to the nestedness issue with a geometrical object that has not been much considered yet in machine learning: flags—i.e. sequences of nested subspaces of increasing dimension. The idea is to first select a sequence of candidate dimensions—the signature of the flag—then formulate a multilevel subspace learning criterion—with the flag trick—and finally optimize over the space of flags of the chosen signature, which has a manifold structure similar to the one of Grassmannians. The output subspaces are naturally nested, which solves the nestedness issue of subspace learning methods and provides the subspaces with a hierarchy that is good for interpretability (Huber, 1985, Section 7). Moreover, the nested representations of the data can be fit to general machine learning algorithms and combined via ensembling methods, in which case the weights can be interpreted as a measure of importance for the different dimensions, echoing the perspectives of the celebrated paper on the automatic choice of dimensionality in PCA (Minka, 2000). Beyond the change of paradigm (from subspaces to flags of subspaces), the main contribution of the paper is the flag trick: a generic method to convert a fixed-dimension subspace criterion into a multilevel flag criterion, without any nestedness issue.
The paper is organized as follows. In section 2, we provide some reminders on flag manifolds and describe a steepest descent optimization algorithm that we will use throughout this work. In section 3, we introduce the flag trick and propose an ensemble learning algorithm to leverage the hierarchical information of the flag into richer machine learning models. In section 4, we show the interest of the flag trick to several subspace learning methods such as robust subspace recovery, trace ratio and spectral clustering. In section 5, we conclude and discuss the limits and perspectives of such a framework. In appendix (Appendix A, Appendix B, Appendix C), we show how the flag trick can be used to develop two advanced optimization algorithms on flag manifolds (namely an iteratively reweighted least squares (IRLS) method for robust subspace recovery and a Newton-Lanczos method for trace ratio problems) that go beyond the limitations of simple algorithms like the steepest descent, and we provide the proofs that are not in the main body.
2 Reminders on flag manifolds
A flag is a sequence of nested linear subspaces of increasing dimension. This section introduces flag manifolds and provides the minimal tools to perform optimization on those spaces. Much more details and properties can be obtained in dedicated papers (Ye et al., 2022).
@@ 2.1 Flags in the scientific literature
By providing a natural parametrization for the eigenspaces of symmetric matrices, flags have long been geometrical objects of interest in the scientific community, with traditional applications in physics (Arnold, 1972) and linear algebra (Ammar and Martin, 1986). Modern computational methods on flag manifolds arose later with the seminal work of Edelman et al. (1998) who provided some optimization algorithms on matrix manifolds such as Grassmannians and Stiefel manifolds. Later, Nishimori et al. (2006) provided an optimization algorithm on flag manifolds for independent subspace analysis (Cardoso, 1998; Hyvärinen and Hoyer, 2000). Recently, flag manifolds played a central role in the investigation of a so-called curse of isotropy in principal component analysis (Szwagier and Pennec, 2024).
Flags as a sequence of nested subspaces were first introduced in the machine learning literature as robust prototypes to represent collections of subspaces of different dimensions (Draper et al., 2014; Mankovich et al., 2022). They were obtained via a sequential construction (where one dimension is added at a time), which can be problematic for greediness reasons (Huber, 1985; Lerman and Maunu, 2018b). Pennec (2018) was the first to show that PCA can be reformulated as an optimization problem on flag manifolds by summing the unexplained variance at different dimensions. This principle was recently applied to several variants of PCA under the name of flagification in Mankovich et al. (2024), showing an important robustness to outliers. Finally, some works compute distance-related quantities on flag-valued datasets, with notable applications in computer vision and hyperspectral imaging (Ma et al., 2021; Nguyen, 2022; Mankovich and Birdal, 2023; Szwagier and Pennec, 2023).
@@ 2.2 Definition and representation of flag manifolds
[…]
The set of flags of signature $(p,q_{1:d})$ is a smooth manifold (Ye et al., 2022), denoted here $\operatorname{Fl}(p,q_{1:d})$. Flag manifolds generalize Grassmannians—since $\leavevmode\nobreak\ {\operatorname{Fl}(p,(q,))=\operatorname{Gr}(p,q)}$—and therefore share many practical properties that are useful for optimization (Edelman et al., 1998). In the following, we will frequently use the following notations: $q_{0}:=0$, $q:=q_{d}$ and $q_{d+1}:=p$.
For computational and numerical reasons, flags are often represented as orthonormal $q$-frames. Those correspond to points on the Stiefel manifold $\operatorname{St}(p,q)=\{U\in\mathbb{R}^{p\times q}\colon U^{\top}U=I_{q}\}$). Let us define sequentially, for $k\in[1,d]$, $U_{k}\in\operatorname{St}(p,q_{k}-q_{k-1})$ such that $[U_{1}|\dots|U_{k}]$ is an orthonormal basis of $\mathcal{S}_{k}$ (this is possible thanks to the nestedness of the subspaces). Then, $U_{1:d}:=[U_{1}|\dots|U_{d}]\in\operatorname{St}(p,q)$ is a representative of the flag $\mathcal{S}_{1:d}$. We call it the Stiefel representation of flags. Such a representation is not unique—contrary to the projection representation defined previously—due to the rotational-invariance of orthonormal bases of subspaces. More precisely, if $U_{1:d}$ is a Stiefel representative of the flag $\mathcal{S}_{1:d}$, then for any set of orthogonal matrices $R_{k}\in\mathcal{O}(q_{k}-q_{k-1})$, the matrix $U^{\prime}_{1:d}:=[U_{1}R_{1}|\dots|U_{d}R_{d}]$ spans the same flag of subspaces $\mathcal{S}_{1:d}$. This provides flag manifolds with a quotient manifold structure (Edelman et al., 1998; Absil et al., 2009; Ye et al., 2022):
@@
For computations, one might have to perform the orthogonal completion of some Stiefel representatives. Let $U_{1:d}:=[U_{1}|\dots|U_{d}]\in\operatorname{St}(p,q)$, then one denotes $U_{d+1}\in\operatorname{St}(p,p-q_{d})$ to be any orthonormal basis such that $U_{1:d+1}:=[U_{1}|\dots|U_{d}|U_{d+1}]\in\mathcal{O}(p)$. Such an orthogonal matrix $U_{1:d+1}$ will be called an orthogonal representative of the flag $\mathcal{S}_{1:d}$. In the following, we may abusively switch from one representation to the other since they represent the same flag.

### v2

@@ Abstract
Many machine learning methods look for low-dimensional representations of the data. The underlying subspace can be estimated by first choosing a dimension $q$ and then optimizing a certain objective function over the space of $q$-dimensional subspaces—the Grassmannian. Trying different $q$ yields in general non-nested subspaces, which raises an important issue of consistency between the data representations. In this paper, we propose a simple and easily implementable principle to enforce nestedness in subspace learning methods. It consists in lifting Grassmannian optimization criteria to flag manifolds—the space of nested subspaces of increasing dimension—via nested projectors. We apply the flag trick to several classical machine learning methods and show that it successfully addresses the nestedness issue.
Keywords: subspace learning, Grassmann manifolds, flag manifolds, nested subspaces, dimensionality reduction
@@ 1 Introduction
Finding low-dimensional representations of datasets is a quite common objective in machine learning, notably in virtue of the curse of dimensionality (bellman_dynamic_1984). A classical way of building such a representation is by searching for a low-dimensional subspace that well represents the data, where “well” is defined by an application-dependent criterion (cunningham_linear_2015). Here is a non-exhaustive list of examples of subspace learning problems. Principal component analysis (PCA) (pearson_lines_1901; hotelling_analysis_1933; jolliffe_principal_2002) searches for a low-dimensional subspace that minimizes the average squared Euclidean distance to the data. Robust subspace recovery (RSR) (lerman_overview_2018) minimizes the average (absolute) Euclidean distance to the data, which is less sensitive to outliers. Many matrix decomposition methods, like robust PCA (candes_robust_2011) and matrix completion (keshavan_matrix_2010; candes_exact_2012) look for low-rank approximations of the data matrix, which can be decomposed into a product of subspace and coordinate matrices. Trace ratio (TR) (ngo_trace_2012) refers to a wide class of problems that look for subspaces making a tradeoff between desired yet antagonist properties of low-dimensional embeddings. An example is Fisher’s linear discriminant analysis (LDA) (fisher_use_1936) which seeks to maximize the between-class variance while minimizing the within-class variance. Domain adaptation methods learn some domain-invariant subspaces (baktashmotlagh_unsupervised_2013) by minimizing the projected maximum mean discrepancy (MMD) (gretton_kernel_2012) between the source and target distributions. Subspace tracking methods incrementally minimize a distance between the current subspace and the available data (balzano_online_2010). Subspaces can also be estimated not from the data but from their adjacency matrix or graph Laplacian, as done in the celebrated Laplacian eigenmaps (belkin_laplacian_2003) and spectral clustering (ng_spectral_2001). Subspaces can also be estimated beyond Euclidean data, for instance on symmetric positive-definite matrix datasets (harandi_dimensionality_2018). In all the previously cited examples, the search space—the space of all linear subspaces of dimension $q$ embedded in an ambient space of larger dimension pp—is called the Grassmannian, or the Grassmann manifold (bendokat_grassmann_2024) and is denoted $\operatorname{Gr}(p,q)$. Consequently, many machine learning methods can be recast as an optimization problem on Grassmann manifolds.
Remark 1 (Sequential Methods)
Some alternative methods to subspace learning via Grassmannian optimization exist. Notably, a large family of methods—like canonical correlation analysis (CCA) (hotelling_relations_1936; hardoon_canonical_2004), independent component analysis (ICA) (hyvarinen_independent_2000), partial least squares (PLS) (geladi_partial_1986) or projection pursuit (huber_projection_1985; tsakiris_dual_2018)—builds a low-dimensional subspace out of the data sequentially. They first look for the best 1D approximation and then recursively add dimensions one-by-one via deflation schemes or orthogonality constraints (lerman_overview_2018, Section II.C). Those sequential methods however suffer from many issues—from a subspace estimation point of view—due to their greedy nature. Those limitations are well described in cunningham_linear_2015 and lerman_overview_2018. In light of the fundamental differences between those methods and the previously described subspace methods, we decide not to address sequential methods in this paper.
@@
Learning a low-dimensional subspace from the data through Grassmannian optimization requires to choose the dimensionality $q$ a priori. This prerequisite has many important limitations, not only since a wrong choice of dimension might remove all the theoretical guarantees of the method (lerman_overview_2018), but also because the observed dataset might simply not have a favored dimension, above which the added information is worthless. [^1] Many generative models assume that the data lies on a lower dimensional subspace, up to isotropic Gaussian noise (tipping_probabilistic_1999; lerman_robust_2015). However, anisotropy is much more realistic (maunu_well-tempered_2019), which nuances the theoretical guarantees and removes hopes for a clear frontier between signal and noise at the intrinsic dimension. For these reasons, one might be tempted to “try” several dimensions—i.e. run the subspace optimization algorithm with different dimensions and choose the best one a posteriori, with cross-validation or statistical model selection techniques for instance. In addition to being potentially costly, such a heuristic raises an important issue that we later refer to as the nestedness issue: the optimal subspaces at different dimensions are not nested in general. This notably means that the data embeddings at two different dimensions might drastically differ, which is a pitfall for data analysis.
Remark 2 (Importance of Nestedness)
The importance of nestedness has been largely documented in the statistical literature: it yields a hierarchy of representations, improving interpretability, dimension selection, computational efficiency and stability. In geometric statistics, nestedness is considered as a key feature for multilevel analysis (damon_backwards_2014). Many works consequently directly impose the nestedness of the approximation subspaces by construction (e.g., via forward or backward sequential methods), in contrast to subspace-based methods which cannot guarantee nestedness. One can notably cite the geodesic versions of PCA (fletcher_principal_2004; sommer_optimization_2014; huckemann_intrinsic_2010), the principal nested spheres of jung_analysis_2012, the barycentric subspace analysis of pennec_barycentric_2018 and subsequent contributions (eltzner_torus_2018; huckemann_backward_2018; dryden_principal_2019; jaquier_high-dimensional_2020; yang_nested_2021; fan_nested_2022; mankovich_chordal_2023; rabenoro_geometric_2024; mankovich_flag_2025; su_principal_2025). All these works show the interest of nestedness in statistical inference, with theoretical advantages (e.g., sharper confidence intervals) and meaningful applications. It is also worth noticing that the review paper from lerman_overview_2018 emphasizes that the lack of nestedness is a fundamental issue in robust subspace recovery, which is left open for future research.
We illustrate in Figure 1 the nestedness issue on toy datasets related to three important machine learning problems: robust subspace recovery, linear discriminant analysis and sparse spectral clustering (lu_convex_2016; wang_grassmannian_2017).
@@
In this paper, we propose a natural solution to the nestedness issue with a geometrical object that has not been much considered yet in machine learning: flags—i.e. sequences of nested subspaces of increasing dimension. The idea is to first select a sequence of candidate dimensions—the signature of the flag—then formulate a multilevel subspace learning criterion which extends the original one by integrating multiple dimensions simultaneously—with the flag trick—and finally optimize over the space of flags of the chosen signature, which has a manifold structure similar to the one of Grassmannians. The output subspaces are naturally nested, which solves the nestedness issue of subspace learning methods and provides the subspaces with a hierarchy that is good for interpretability (huber_projection_1985, Section 7). Moreover, the nested representations of the data can be fit to general machine learning algorithms and combined via ensembling methods. The ensembling weights can then be interpreted as a measure of importance for the different dimensions, echoing the perspectives of the celebrated paper on the automatic choice of dimensionality in PCA (minka_automatic_2000). Beyond the change of paradigm (from subspaces to flags of subspaces), the main contribution of the paper is the flag trick: a generic method to convert a fixed-dimension subspace criterion into a multilevel flag criterion, without nestedness issue.
The paper is organized as follows. In Section 2, we provide some prerequisites on flag manifolds and describe a steepest descent optimization algorithm that we will use throughout this work. In Section 3, we introduce the flag trick and propose an ensemble learning algorithm to leverage the hierarchical information of the flag into richer machine learning models. In Section 4, we show the interest of the flag trick to several subspace learning methods such as robust subspace recovery, trace ratio and spectral clustering. In Section 5, we conclude and discuss the limits and perspectives of such a framework. In Appendix A, B, C, we show how the flag trick can be used to develop two advanced optimization algorithms on flag manifolds (namely an iteratively reweighted least squares (IRLS) method for robust subspace recovery and a Newton-Lanczos method for trace ratio problems) that go beyond the limitations of simple algorithms like the steepest descent, and we provide the proofs that are not in the main body. In Appendix D, we report the empirical running times for the flag trick in robust subspace recovery, trace ratio and spectral clustering.
2 Basics of Flag Manifolds
A flag is a sequence of nested linear subspaces of increasing dimension. This section introduces flag manifolds and provides the minimal tools to perform optimization on those spaces. More details and properties can be obtained in dedicated papers (ye_optimization_2022).
@@ 2.1 Flags in the Scientific Literature
By providing a natural parametrization for the eigenspaces of symmetric matrices, flags have long been geometrical objects of interest in the scientific community, with traditional applications in physics (arnold_modes_1972) and numerical analysis (duistermaat_functions_1983; ammar_geometry_1986; watkins_chasing_1991; helmke_isospectral_1991; helmke_optimization_1994). The seminal work of edelman_geometry_1998 on the geometry of algorithms with orthogonality constraints paved the way for Riemannian optimization algorithms on Stiefel, Grassmann and related matrix manifolds. The first work that explicitly proposed an optimization algorithm on flag manifolds dates back to 2006. nishimori_riemannian_2006 reformulated independent subspace analysis—an important variant of independent component analysis which maximizes the independence between the norms of projected samples into mutually-orthogonal subspaces (cardoso_multidimensional_1998; hyvarinen_emergence_2000)—as an optimization problem on flag manifolds, and solved it via a Riemannian gradient descent. This led to many papers developing or performing optimization on flag manifolds (nishimori_flag_2007; nishimori_natural_2008; fiori_extended_2011; fiori_extended_2012; ye_optimization_2022; zhu_practical_2024) and to implementations in open-source packages for differential geometry such as Manifolds.jl (axen_manifoldsjl_2023).
Recently, flag manifolds played a central role in the investigation of a curse of isotropy in principal component analysis. By means of a probabilistic covariance model parameterized with flags, szwagier_curse_2025 show that close eigenvalues—whose relative distance is lower than a certain threshold—should be assumed equal. The resulting principal subspace analysis has led to several follow-up works, notably an efficient algorithm for covariance estimation in high dimensions (szwagier_eigengap_2025) and an extension to parsimonious Gaussian mixture models (szwagier_parsimonious_2025).
It is important to notice that flags—as mutually-orthogonal subspaces—also implicitly appear in many works, with different keywords. For instance, the mutually-orthogonal class-subspaces of watanabe_subspace_1973, the adaptive-subspace self-organizing maps of kohonen_emergence_1996 and follow-up works (tae-kyun_kim_-line_2010; giguere_manifold_2017) implicitly involve flags under the name of “mutually-orthogonal subspaces”. Flags also appear in many spectral methods under keywords such as “isospectral manifolds” (deift_ordinary_1983; watkins_isospectral_1984; brockett_dynamical_1991; helmke_isospectral_1991; lim_simple_2024); indeed, at fixed eigenvalues, the remaining degrees of freedom in a symmetric matrix are the associated eigenspaces, which exactly correspond to flags.
Flags as sequences of nested subspaces were first introduced in the machine learning literature as robust prototypes to represent collections of subspaces of different dimensions (draper_flag_2014; santamaria_order_2016; mankovich_flag_2022). They were obtained via a sequential construction (where one dimension is added at a time), which can be problematic for greediness reasons (huber_projection_1985; lerman_overview_2018). pennec_barycentric_2018 was the first to show that PCA can be reformulated as an optimization problem on flag manifolds by summing the unexplained variance at different dimensions, raising interesting perspectives for multilevel data analysis on manifolds. This principle was recently applied to several variants of PCA under the name of flagification in mankovich_fun_2024, showing an important robustness to outliers. Some works compute distance-related quantities on flag-valued datasets, with notable applications in computer vision and hyperspectral imaging (ma_flag_2021; ma_self-organizing_2022; nguyen_closed-form_2022; mankovich_chordal_2023; szwagier_rethinking_2023). Other applications include shape spaces (ciuclea_shape_2023), quantum physics and chemistry (vidal_geometric_2024). Finally, flags were recently used to learn nested representations from data with a prespecified hierarchy (mankovich_flag_2025).
@@ 2.2 Definition and Representation of Flag Manifolds
Let $p\geq 2$ and $q_{1:d}\coloneqq(q_{1},q_{2},\dots,q_{d})$ be a sequence of increasing integers such that $0<q_{1}<q_{2}<\cdots<q_{d}<p$. A flag of signature $(p,q_{1:d})$ is a sequence of nested linear subspaces $\{0\}\subset\mathcal{S}_{1}\subset\mathcal{S}_{2}\subset\dots\subset\mathcal{S}_{d}\subset\mathbb{R}^{p}$ of respective dimensions $q_{1},q_{2},\dots,q_{d}$, noted here $\mathcal{S}_{1:d}\coloneqq(\mathcal{S}_{1},\dots,\mathcal{S}_{d})$. [^2] Flags can equivalently be defined as sequences of mutually orthogonal subspaces $\mathcal{V}_{1}\perp\mathcal{V}_{2}\dots\perp\mathcal{V}_{d}$, of respective dimensions $q_{1},q_{2}-q_{1},\dots,q_{d}-q_{d-1}$, by taking $\mathcal{V}_{k}$ to be the orthonormal complement of $\mathcal{S}_{k-1}$ onto $\mathcal{S}_{k}$. This definition is more convenient for computations, but we won’t need it in this paper. A flag $\mathcal{S}_{1:d}$ can be canonically represented as a sequence of symmetric matrices that are the orthogonal projection matrices onto the nested subspaces, i.e. $\mathcal{S}_{1:d}\cong(\Pi_{\mathcal{S}_{1}},\dots,\Pi_{\mathcal{S}_{d}})\in\operatorname{Sym}(p)^{d}$. We call it the projection representation of flags.
The set of flags of signature $(p,q_{1:d})$ is a smooth manifold (ye_optimization_2022), denoted here $\operatorname{Fl}(p,q_{1:d})$. Flag manifolds generalize Grassmannians when d=1d=1—since $\penalty 10000\ {\operatorname{Fl}(p,q)=\operatorname{Gr}(p,q)}$—and therefore share many practical properties that are useful for optimization (edelman_geometry_1998). In the following, we will frequently use the following notations: $_{0}\coloneqq 0, qq$$\coloneqq q_{d}$ and $q_{d+1}\coloneqq p$.
For computational and numerical reasons, flags are often represented as orthonormal $q$-frames. Those correspond to points on the Stiefel manifold $\operatorname{St}(p,q)=\{U\in\mathbb{R}^{p\times q}\colon U^{\top}U=I_{q}\}$). Let us define sequentially, for $k\in\{1,\dots,d\}$, $U_{k}\in\operatorname{St}(p,q_{k}-q_{k-1})$ such that $\begin{bmatrix}U_{1}&\cdots&U_{k}\end{bmatrix}$ is an orthonormal basis of $\mathcal{S}_{k}$ (this is possible thanks to the nestedness of the subspaces). Then, $U_{1:d}\coloneqq\begin{bmatrix}U_{1}&\cdots&U_{d}\end{bmatrix}\in\operatorname{St}(p,q)$ is a representative of the flag $\mathcal{S}_{1:d}$. We call it the Stiefel representation of flags. Such a representation is not unique—contrary to the projection representation defined previously—due to the rotational-invariance of orthonormal bases of subspaces. More precisely, if $U_{1:d}$ is a Stiefel representative of the flag $\mathcal{S}_{1:d}$, then for any set of orthogonal matrices $R_{k}\in\mathcal{O}(q_{k}-q_{k-1})$, the matrix $U^{\prime}_{1:d}\coloneqq\begin{bmatrix}U_{1}R_{1}&\cdots&U_{d}R_{d}\end{bmatrix}$ spans the same flag of subspaces $\mathcal{S}_{1:d}$. This provides flag manifolds with a quotient manifold structure (edelman_geometry_1998; absil_optimization_2008; ye_optimization_2022):
@@
For computations, one might have to perform the orthogonal completion of some Stiefel representatives. Let $U_{1:d}\coloneqq\begin{bmatrix}U_{1}&\cdots&U_{d}\end{bmatrix}\in\operatorname{St}(p,q)$, then one denotes $U_{d+1}\in\operatorname{St}(p,p-q_{d})$ to be any orthonormal basis such that $U_{1:d+1}\coloneqq\begin{bmatrix}U_{1}&\cdots&U_{d}&U_{d+1}\end{bmatrix}\in\mathcal{O}(p)$. Such an orthogonal matrix $U_{1:d+1}$ will be called an orthogonal representative of the flag $\mathcal{S}_{1:d}$. In the following, we may abusively switch from one representation to the other since they represent the same flag.